OpenStudy (anonymous):

what is the point of discontinuity of f(x) = x^2-64/x-8 Could someone please explain this to me? Thanks.

hero (hero):

Let me see if I can. Basically, the question is since \[\frac{x^2 - 64}{x - 8} = x + 8\] However they are only equal if \(x \ne 8\)

OpenStudy (anonymous):

if denominator becomes zero, the function will become undefined so the discontinuity of the function will be the zero of the denominator.

hero (hero):

In other words, if you graphed \[\frac{x^2 - 64}{x - 8}\] There would be an empty circle in the graph at x = 8

OpenStudy (anonymous):

so that would be considered the point of discontinuity?

OpenStudy (anonymous):

Or would it be 8,16?

OpenStudy (anonymous):

however,if you factor out the numerator the denominator will cancel out, so the function will have a hole at x=8...

OpenStudy (anonymous):

so the point of disconitunity is where there is a hole?

OpenStudy (anonymous):

and that hole is at 8?

OpenStudy (anonymous):

yes you are right..

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

no problem